# CLIMB: A Benchmark of Clinical Bias in Large Language Models

### v1

@@ 4.3.2 Sex-specific Samples
To analyze the breakdown contributor of the overall performance differences of sex, we divide our dataset into sex-specific and sex-neutral sample groups according to the criteria introduced in Section 3. In Table 5, we separately report their extrinsic bias scores when only replacing sex. In sex-neutral samples, performance differences (i.e., sex bias) occur for all models. In sex-specific samples, several models (# 7, 9 and 13) show counter-intuitive performance increase. For all medically adapted models, their performance differences are even larger than their based counterparts in sex-neutral samples (i.e., more biased), which suggests that their medically adapted training may not help to identify these samples are sex-neutral, therefore unable to avoid sex bias when predicting diagnoses.
@@ 5 Conclusion
In conclusion, CLIMB serves as a pioneering benchmark specifically crafted to concurrently and systematically evaluate both intrinsic and extrinsic bias in LLMs applied for clinical decision-making tasks. Notably, we introduce a novel metric, AccocMAD, allowing for a more comprehensive assessment of disparities across multiple demographic groups, and several advanced strategies to adapt counterfactual intervention into the evaluation of extrinsic bias in clinical setting. Our extensive experiments across a range of popular and medically adapted LLM models, particularly from the Mistral and LLaMA families, unveil prevalent biased behaviors, with even some medically adapted LLMs performing worse than their base counterparts and lacking effective bias mitigation. This work not only underscores the critical need for bias mitigation in medically adapted LLM applications but also sets a new standard for future bias evaluations of AI systems in healthcare.
@@
Appendix A Evaluation Set Statistics
Figure 2 and Figure 3 show the statistics in CLIMB evaluation set. Comparing subgraphs (a) and (b), it is evident that our subset maintains the sample diversity of the CliBench dataset.
Figure 2: Admission Distribution Comparison
Figure 3: Diagnosis Distribution Comparison

### v2

@@ 4.3.2 Sex-specific Samples
To analyze the breakdown contributor of the overall performance differences of sex, we divide our dataset into sex-specific and sex-neutral sample groups according to the criteria introduced in Section 3. In Table 5, we separately report their extrinsic bias scores when only replacing sex. In sex-neutral samples, performance differences (i.e., sex bias) occur for all models. In sex-specific samples, several models (# 7, 9 and 13) show counter-intuitive performance increase. We observe this is because the majority of sex-neutral diagnoses within these samples contribute more positively than the negative impact of the male-specific diagnoses. However, the positive changes of sex-neutral diagnoses are still biased as we expect no change for them when perturbing sex. For all medically adapted models, their performance differences are even larger than their based counterparts in sex-neutral samples (i.e., more biased), which suggests that their medically adapted training may not help to identify these samples are sex-neutral, therefore unable to avoid sex bias when predicting diagnoses.
@@ 5 Conclusion
In conclusion, CLIMB serves as a pioneering benchmark specifically crafted to concurrently and systematically evaluate both intrinsic and extrinsic bias in LLMs applied for clinical decision-making tasks. Notably, we introduce a novel metric, AccocMAD, allowing for a more comprehensive assessment of disparities across multiple demographic groups, and several advanced strategies to adapt counterfactual intervention into the evaluation of extrinsic bias in clinical setting. Our extensive experiments across a range of popular and medically adapted LLM models, particularly from the Mistral and LLaMA families, unveil prevalent biased behaviors, with even some medical adapted LLMs performing worse than their base counterparts and lacking effective bias mitigation. This work not only underscores the critical need for bias mitigation in medical adapted LLM applications but also sets a new standard for future bias evaluations of AI systems in healthcare.
@@
Appendix A Prompt Template and Evaluation Set
Table 6 show the our prompt template for extrinsic bias evaluation which is the same as CliBench. This prompt of diagnsis task is realistical for clinical bias evaluation because 1) It’s realistically used by clinical resource management units for resource forecasting, billing departments for auditing and insurance companies for risk reduction; 2) Infering diagnosis could indirectly benefit patient treatment, such as medication recommendation for online pharmacy and ordering lab tests; 3) It examines a comprehensive set of clinical capabilities of the LM covering a long chain of clinical operations, including clinical note understadning, diagnosis decision space understanding, reasoning with complex information, and clinical knowledge about facts about lab test and drugs, and associations between sympothms and diagnosis.
System Prompt You are a professional clinician in a hospital with expert knowledge in medical and clinical domains. The task is to make a list of diagnoses for this patient based on the provided information of the patient. The diagnosis can be in ICD-10-CM code format (such as S12.000G), or natural language description of the disease. Separate each diagnosis with a new line. Please provide as many diagnoses as you can until you are not confident about your diagnosis decision. User Prompt Patient information: sex is {}, ethnicity is {} … {Clinical Note … } What are the diagnoses for this patient?
Table 6: Prompt template of the diagnosis task in CliBench
Figure 2 and Figure 3 show the statistics of CLIMB evaluation set. Comparing sub-figures (a) and (b), it is evident that our subset maintains the sample diversity of the CliBench dataset.
Figure 2: Admission distribution comparison.
Figure 3: Diagnosis distribution comparison.
@@ Appendix B Original Performance of Evaluation Set for Extrinsic Bias
Appendix C Name Resampling for Intrinsic Bias
Table 8 shows the results of different top-K name resampling. We observe that our key observations from Table 1 ($k=5$) remain consistent across all resamplings.
# Model/Method Params L1 L2 L3 L4 L5 Avg
$k=1$
Mistral Instruct v0.1 7B 0.35 0.41 0.49 0.69 0.92 0.57
Mistral Instruct v0.2 7B 0.33 0.39 0.48 0.65 0.84 0.54
Mistral Instruct v0.3 7B 0.57 0.68 0.72 0.83 1.01 0.76
BioMistral DARE 7B 0.40 0.44 0.53 0.72 0.92 0.60
Mistral Instruct v0.1 46.7B 0.73 0.83 0.97 1.08 1.24 0.97
LLaMA2 Chat 7B 0.74 0.75 0.77 0.77 0.79 0.76
Meditron 7B 0.22 0.27 0.34 0.40 0.50 0.35
Asclepius 7B 0.97 0.95 0.95 0.98 0.98 0.97
LLaMA2 Chat 13B 1.11 1.10 1.10 1.08 1.13 1.10
LLaMA2 Chat 70B 0.81 0.83 0.84 0.88 0.87 0.85
LLaMA3 Instruct 8B 0.53 0.55 0.56 0.58 0.61 0.57
OpenBioLLM 8B 0.96 0.95 0.93 0.94 0.99 0.95
LLaMA3 Instruct 70B 0.88 0.89 0.88 0.84 0.83 0.86
$k=2$
Mistral Instruct v0.1 7B 0.24 0.29 0.36 0.53 0.74 0.43
Mistral Instruct v0.2 7B 0.23 0.27 0.34 0.47 0.62 0.39
Mistral Instruct v0.3 7B 0.38 0.50 0.56 0.64 0.82 0.58
BioMistral DARE 7B 0.28 0.32 0.39 0.56 0.74 0.46
Mistral Instruct v0.1 46.7B 0.54 0.66 0.80 0.90 1.04 0.79
LLaMA2 Chat 7B 0.50 0.53 0.56 0.56 0.59 0.55
Meditron 7B 0.17 0.20 0.26 0.30 0.39 0.26
Asclepius 7B 0.78 0.76 0.75 0.77 0.78 0.77
LLaMA2 Chat 13B 0.93 0.93 0.96 0.95 1.02 0.96
LLaMA2 Chat 70B 0.56 0.59 0.62 0.64 0.65 0.61
LLaMA3 Instruct 8B 0.56 0.56 0.56 0.57 0.57 0.56
OpenBioLLM 8B 0.73 0.74 0.73 0.72 0.78 0.74
LLaMA3 Instruct 70B 0.73 0.75 0.75 0.71 0.70 0.73
$k=3$
Mistral Instruct v0.1 7B 0.18 0.23 0.29 0.44 0.64 0.36
Mistral Instruct v0.2 7B 0.16 0.20 0.26 0.36 0.48 0.29
Mistral Instruct v0.3 7B 0.29 0.42 0.48 0.55 0.70 0.49
BioMistral DARE 7B 0.21 0.24 0.31 0.47 0.65 0.38
Mistral Instruct v0.1 46.7B 0.43 0.55 0.68 0.79 0.93 0.68
LLaMA2 Chat 7B 0.84 0.84 0.84 0.89 0.92 0.87
Meditron 7B 0.14 0.18 0.22 0.26 0.34 0.23
Asclepius 7B 0.76 0.74 0.73 0.71 0.70 0.73
LLaMA2 Chat 13B 0.78 0.77 0.81 0.84 0.96 0.83
LLaMA2 Chat 70B 0.39 0.44 0.47 0.50 0.54 0.47
LLaMA3 Instruct 8B 0.42 0.44 0.45 0.45 0.45 0.44
OpenBioLLM 8B 0.61 0.62 0.61 0.59 0.64 0.61
LLaMA3 Instruct 70B 0.58 0.60 0.61 0.61 0.61 0.60
$k=4$
Mistral Instruct v0.1 7B 0.16 0.20 0.25 0.39 0.59 0.32
Mistral Instruct v0.2 7B 0.14 0.17 0.22 0.30 0.42 0.25
Mistral Instruct v0.3 7B 0.27 0.38 0.43 0.48 0.62 0.44
BioMistral DARE 7B 0.19 0.21 0.28 0.43 0.61 0.34
Mistral Instruct v0.1 46.7B 0.35 0.48 0.60 0.70 0.86 0.60
LLaMA2 Chat 7B 0.78 0.79 0.80 0.85 0.88 0.82
Meditron 7B 0.13 0.16 0.20 0.24 0.32 0.21
Asclepius 7B 0.73 0.71 0.70 0.68 0.67 0.70
LLaMA2 Chat 13B 0.90 0.93 0.98 1.01 1.10 0.98
LLaMA2 Chat 70B 0.39 0.43 0.47 0.51 0.54 0.47
LLaMA3 Instruct 8B 0.37 0.39 0.40 0.40 0.40 0.39
OpenBioLLM 8B 0.55 0.56 0.56 0.54 0.59 0.56
LLaMA3 Instruct 70B 0.53 0.55 0.56 0.58 0.58 0.56
Table 8: Results of top-K name resampling for overall intrinsic bias.